OpenStudy (anonymous):

how to show your work for |n-4|=9

OpenStudy (amistre64):

doesnt this mean that: n-4 = 9 or n-4 = -9

OpenStudy (amistre64):



OpenStudy (anonymous):

n-4=9 add 4 to both sides to eliminate n-4+4=9+4 n=13

OpenStudy (anonymous):

i did not see your I I parentheses, sorry, do not use my answer.